A parallel plate capacitor with plate separation of 4.0 cm has a plate area of 4.0 × 10-2 m2. What is the capacitance of this capacitor with air between theseplates?

a) 8.9 × 10-11 F
b) 8.9 × 10-15 F
c) 8.9 × 10-12 F
d) 8.9 × 10-14 F
e) 8.9 × 10-13 F
Answer #1
Capacitance = A*Epsilon_0/d
= 4.0 × 10-2 * 9*10^-12/.04 = 4*9/.04 *10^10 = 900 * 10^-14
= 9 pF
answered by: Lolerskater
Answer #2
C = 8.85*10^-12*4*10^-2/0.04 =8.9 *10^-12 F =8.9 pF

option C
answered by: Nandi
Answer #3
C=Ae0/d=4x10^-2x/(4x10^-2)
C= 8.9 × 10-12 F
answered by: jim bean
Answer #4
8.9 × 10-12 F
answered by: zohra